# Exact Simulation of Longitudinal Data from Marginal Structural Models

### v2

@@ Abstract
Simulating longitudinal data from specified marginal structural models is a crucial but challenging task for evaluating causal inference methods and clinical trial design. While data generation typically proceeds in a fully conditional manner using structural equations according to a temporal ordering, it is difficult to ensure alignment between conditional distributions and the target marginal causal effects. This misalignment presents a fundamental challenge in simulating data that adheres to marginal structural model specifications. To address this, we propose a flexible and efficient algorithm for simulating longitudinal data that adheres exactly to a specified marginal structural model. Recognizing the importance of time-to-event outcomes in clinical research, we extend our approach to accommodate survival models. Compared to existing approaches, our method offers several advantages: it enables exact simulation from a known causal model rather than relying on approximations; avoids restrictive assumptions about the data-generating process; and remains computationally efficient by requiring only the evaluation of analytic functions. This last benefit contrasts with methods that use computationally intensive techniques such as Monte Carlo approximations or numerical integration. Through simulation studies replicating realistic scenarios, we validate the method’s accuracy and utility. Our method will facilitate researchers in effectively simulating data with target causal structures for their specific scenarios.
@@ 1 Introduction
Longitudinal data are fundamental for studies analyzing the causal effect of a treatment or exposure on an outcome, especially for research into chronic diseases such as diabetes or HIV. Unlike cross-sectional studies, longitudinal designs capture the dynamic nature of treatment regimens, which often vary over time due to clinical decisions, patient adherence, or external factors. Even in clinical trials with strict adherence protocols, intercurrent events—such as treatment discontinuation or supplementary medication use—are common. By accounting for these treatment dynamics, longitudinal data enable the estimation of meaningful causal quantities, such as the “on-treatment” estimand, which quantifies the treatment effect assuming treatment is administered as intended.
@@
A central challenge in analyzing longitudinal data is addressing time-varying confounding, where a time-dependent covariate predicts both the outcome and future treatments, while itself being influenced by past treatments. For example, in the case study in Robins et al. (2000) on the effect of zidovudine (AZT) treatment on HIV infection, the biomarker CD4 lymphocyte count is both affected by prior AZT treatment and is predictive of future therapy initiation and HIV progression. The presence of such time-varying confounding renders conventional adjustment methods, like regression and stratification, generally biased.
@@
Marginal structural models, introduced by Robins (1997), have emerged as a powerful tool for estimating causal effects under time-varying confounding. Model parameters can be consistently estimated using inverse probability of treatment weighting (Robins, 1999, 2000). Among alternative methods, including structural nested models estimated through g-computation (Robins, 1989; Robins and Tsiatis, 1991), marginal structural models have gained prominence due to their interpretability and ease of implementation. A systematic review by Clare et al. (2019) highlights their widespread use in applied research.
@@
As causal inference methods for longitudinal data are continually proposed to address complex real-world settings, simulation studies are essential to evaluate their effectiveness for practical use. These studies assess important properties such as bias, confidence interval coverage, relative efficiency, and robustness to violations of the assumptions used, guiding both methodological development and application. Sample size calculations also benefit from simulations. Prior to conducting a trial or an observational study, it is crucial to determine the number of participants or number of observational units required to achieve adequate statistical power. While standard formulas can be used for time-fixed treatments, more complex scenarios with time-varying treatments make deriving analytical formulas impractical or impossible. In such cases, simulation is invaluable, allowing researchers to estimate sample sizes based on realistic data-generating processes and causal structures.
@@
A fundamental requirement for both method evaluation and sample size determination is the ability to simulate data that conform to a specified causal model. However, simulating longitudinal data from a pre-specified marginal structural model is not a straightforward task. While generating data sequentially from structural equations—creating covariates, treatment status, and then the outcome conditioned on past information—may seem intuitive, these models rely on conditional rather than marginal specifications. This makes it difficult to sequentially simulate with conditional distributions in a way that aligns with a pre-specified marginal model; in particular, it is extremely challenging to sample from a null hypothesis —the so-called g-null paradox (Robins and Wasserman, 1997).
@@
The remainder of this paper is structured as follows: Section 2 reviews existing methods for simulating longitudinal data from causal models. Section 3 introduces key preliminaries, including the frugal parameterization (Evans and Didelez, 2024) and copulas. In Section 4, we introduce our proposed methodology and illustrate it in Section 4.4 through a simulation study replicating a real-world study in epidemiology (Bodnar et al., 2004). Section 5 extends the simulation approach to time-to-event data, enabling simulations from survival models such as Cox marginal structural models, including an illustrative example. The paper concludes with a discussion in Section 6. Example code for the simulations in Sections 4.4 and 5.3 is available at https://github.com/XiLinStats/LongSiMSM.
@@ 2 Related Literature
    Simulating from conditional models and deriving or approximating the corresponding marginal models post-hoc.
@@
Several studies have explored the first approach. Xiao et al. (2010) used sequential conditional models to approximate Cox marginal structural models under the rare disease assumption, i.e. very low failure rates, which mitigates the impact of noncollapsibility but assumes the absence of unmeasured confounding between time-varying confounders and outcomes. Young and Tchetgen Tchetgen (2014) employed the g-formula to bridge conditional and marginal models in discrete-time Cox marginal structural models. While theoretically feasible for discrete time-varying covariates, deriving marginal models becomes intractable for continuous covariates due to the need for numerical integration. Simplifications, such as the Markov property or rare disease assumptions, are often necessary but limit generalizability. Sofrygin et al. (2017) developed the simcausal R package for flexible simulation from non-parametric structural equation models. The workflow involves specifying structural equations and sequentially simulating data from conditional distributions. However, deriving analytic expressions for marginal causal quantities under this conditional parameterization is often intractable and prone to model misspecification. To address this, users typically specify low-dimensional working marginal structural models (Neugebauer and van der Laan, 2007) and estimate their parameters via Monte Carlo simulations of counterfactual data. This introduces two limitations: first, the true marginal model remains unknown; second, reverse-engineering simulations to match a target causal function is challenging. Keogh et al. (2021) described simulation from additive-hazard marginal structural models (Aalen, 1989). They utilized the result that if the conditional hazard model follows an additive form, the equivalent marginal hazard model will also be additive. While the marginal model retains the correct form, true parameter values are generally intractable and require Monte Carlo approximations. A further limitation is the difficulty in ensuring non-negative hazards for continuous covariates.
@@
Alternative methods focus on simulating directly from the target marginal model. Havercroft and Didelez (2012) simulated directly from a discrete Cox marginal structural model, introducing time-varying confounding indirectly via a shared latent variable. However, this approach assumes no direct effect of time-varying confounders on outcomes, which may not hold in practice, e.g. the effect of CD4 count on HIV survival (Robins et al., 2000). Evans and Didelez (2024) proposed a frugal parameterization that prioritizes the causal margin and parameterizes the surrounding variables accordingly, enabling direct simulation from the exact marginal model while explicitly modeling time-varying covariates. This addresses key limitations of earlier methods. Building on this, Seaman and Keogh (2024) applied the frugal parameterization to simulate survival outcomes, using a one-dimensional “risk score” to summarize multi-dimensional covariates. While simplifying dependence structures, this method relies on computationally expensive Monte Carlo approximations to estimate the cumulative distribution functions of the risk score, making it infeasible for continuous treatments or baseline variables.
@@
The aims for a new simulation approach therefore include the ability to: precisely specify the marginal structural model; include multiple time-varying covariates, treatments, outcomes, and censoring; have good control over their distributions; avoid restrictive assumptions about the data-generating process. In addition, we would like it to scale well in the number of time-points and variables, and be applicable to real-world problems.
@@ 3.2 Copulas
We can think of continuous multivariate distributions as the joining together of the univariate marginals (or margins) and the dependence structure. This structure can be represented by a copula, which is a multivariate distribution function $C:[0,1]^{d}\rightarrow[0,1]$ with margins uniformly distributed on $(0,1)$. According to Sklar’s theorem (Sklar, 1959, 1973), any continuous multivariate distribution $F$ with marginals $F_{1},\ldots,F_{d}$ can be expressed as
@@
where $\bm{X}=(X_{1},\ldots,X_{d})$ is a vector of random variables, and $u_{i}=F_{i}(x_{i})$, $i=1,\ldots,d$. The copula then takes the form
@@
The choice of copula can be considered independently of the univariate margins.
@@ 3.2.1 Pair-copula construction
Following the general formula given in Joe (1997), the conditional distribution of one variable $y$ given $d$ other variables can be decomposed to $d$ bivariate copula density functions, each of which can be specified flexibly:
@@ 3.2.2 Simulating from pair-copula constructions
The pair-copula construction includes conditional distributions of the form $F(y\,|\,\bm{x})$. Joe (1996) demonstrated the following result:
$F(y\,|\,\bm{x})=\frac{\partial C_{YX_{j}\mid\bm{X}_{-j}}\left\{F(y\,|\,\bm{x}_{-j}),F(x_{j}\,|\,\bm{x}_{-j})\right\}}{\partial F(x_{j}\,|\,\bm{x}_{-j})}.$ (3)
For notational convenience, we introduce the h-function $h(u_{1},u_{2};\Theta)$ from Aas et al. (2009). Let $u_{1}$ and $u_{2}$ be random variables with a uniform distribution on $(0,1)$. The $h$-function is defined as
@@
where the second argument in $h(\cdot;\Theta)$ is the conditional variable, and $\Theta$ represents the set of copula parameters. The $h$-function is effectively the conditional copula $C(u_{1}\,|\,u_{2};\Theta)$.
@@
The derivation of the analytical form of $h(\cdot)$ for common bivariate copulas is straightforward by definition for parametric families. Examples include Gaussian, Student t, Clayton, Frank, and Gumbel copulas.
@@ 4 Simulation Algorithm
We first consider a longitudinal study with $n$ individuals observed at discrete time points, $\mathcal{K}=\{0,\ldots,K\}$. At each follow-up $k$, treatment status $A_{k}$, and a set of time-varying covariates $L_{k}$ are recorded. Additionally, $Z$ represents a set of static baseline covariatesatent variables, if present, are written as $\{H_{k}\}_{k\in\mathcal{K}}$. To indicate the history of a variable up to time $k$, we use bar notation. For instance, $\bar{A}_{k}=\{A_{0},\ldots,A_{k}\}$ and $\bar{L}_{k}=\{L_{0},\ldots,L_{k}\}$, with the convention $\bar{A}_{-1}=\emptyset$ and $\bar{L}_{-1}=\emptyset$. The outcome $Y$ is observed at the final time point $K$. Figure 1 is a graphical representation of this setup, which is common in medical and epidemiological research. Examples include examining the effects of an HIV treatment on infection status assessed at the end of the study period (Robins et al., 2000), and assessing the impact of iron supplement use during pregnancy on the odds of anemia at delivery (Bodnar et al., 2004). In Section 4.4, we illustrate our method through a simulation study that replicates such settings.
@@ 4.2 Frugal parameterization for outcome simulation
$F(y\,|\,\bar{l}_{K},\bar{a}_{K})$ $=h\big{[}\dots h\{h(F^{*}(y\,|\,do(\bar{a}_{K})),F(l_{0});\Theta_{0}),F(l_{1}\,|\,l_{0},a_{0});\Theta_{1}\}\dots;\Theta_{K}\big{]},$
which is centered around the causal margin $F^{*}(y\,|\,\bar{a}_{K})$.
Through inversion the outcome can be simulated in two steps:
@@
    $u^{*}_{Y}$ $=h^{-1}\big{[}\dots h^{-1}\big{\{}h^{-1}(w_{Y},F(l_{K}\,|\,\bar{l}_{K-1},\bar{a}_{K-1});\Theta_{K}),$
    $\qquad\qquad\qquad F(l_{K-1}\,|\,\bar{l}_{K-2},\bar{a}_{K-2});\Theta_{K-1}\big{\}}\dots,F(l_{0});\Theta_{0}\big{]}$ (8)
    $=h^{-1}\big{[}\dots h^{-1}\big{\{}h^{-1}(w_{y},w_{K,1};\Theta_{K}),w_{K-1,1};\Theta_{K-1}\big{\}},\ldots,w_{0,1};\Theta_{0}\big{]},$
@@
An intuitive way to understand the calculation in Step 1 is that the inverse $h$-functions “disentangle” the dependencies between $Y$ and each of $L_{K},L_{K-1},\dots L_{0}$ sequentially, and remove the conditioning of $Y$ on $\bar{L}_{K}$ until the marginal quantile of $Y$ is obtained. The quantiles in the inverse $h$-functions, such as $F(l_{k}\,|\,\bar{l}_{k-1},\bar{a}_{k-1})$, correspond exactly to the random standard uniform variables $w_{k,1}$ used to simulate $l_{k}$ at time step $k$. This enables direct substitution of the $l_{k}$ quantiles into (8) without additional computation.

### v3

@@ Abstract
Simulating longitudinal data from specified marginal structural models is a crucial but challenging task for evaluating causal inference methods and informing study design. While data generation typically proceeds in a fully conditional manner using structural equations according to a temporal ordering, it is difficult to ensure alignment between conditional distributions and the target marginal causal effects, which presents a fundamental challenge. To address this, we propose a flexible and efficient algorithm for simulating longitudinal data that adheres exactly to a specified marginal structural model. Our approach accommodates time-to-event outcomes and extends naturally to survival settings, which are prevalent in applied research. Compared to existing approaches, it offers several advantages: it enables exact simulation from a known causal model rather than relying on approximations; avoids restrictive assumptions about the data-generating process; and remains computationally efficient by requiring only the evaluation of analytical expressions, rather than Monte Carlo methods or numerical integration. Through simulation studies replicating realistic scenarios, we validate the method’s accuracy and utility. Our method will facilitate researchers in effectively simulating data with target causal structures for their specific scenarios.
@@ 1 Introduction
Longitudinal data are fundamental for studies analyzing the causal effect of a treatment or exposure, especially for research into chronic diseases such as diabetes or HIV. Unlike cross-sectional studies, longitudinal designs capture the dynamic nature of treatment regimens, which often vary over time due to clinical decisions, patient adherence, or external factors. Even in clinical trials with strict adherence protocols, intercurrent events—such as treatment discontinuation or supplementary medication use—are common. By accounting for these treatment dynamics, longitudinal data enable the estimation of meaningful causal quantities, such as the “on-treatment” estimand, which quantifies the treatment effect assuming treatment is administered as intended.
@@
A central challenge in analyzing longitudinal data is addressing time-varying confounding, where a time-dependent covariate predicts both the outcome and future treatments, while itself being influenced by past treatments. For example, in the case study by Robins et al. (2000) on the effect of zidovudine (AZT) treatment on HIV infection, the biomarker CD4 lymphocyte count is both affected by prior AZT treatment and is predictive of future therapy initiation and HIV progression. The presence of such time-varying confounding renders conventional adjustment methods, like regression and stratification, generally biased.
@@
Marginal structural models, introduced by Robins (1997), have emerged as a powerful tool for estimating causal effects under time-varying confounding. Model parameters can be consistently estimated using inverse probability of treatment weighting (Robins, 1999, 2000). Among alternative methods, including structural nested models estimated through g-estimation (Robins, 1989, 1992; Robins and Tsiatis, 1991), marginal structural models have gained prominence due to their interpretability and ease of implementation. A systematic review by Clare et al. (2019) highlights their widespread use in applied research.
@@
As causal inference methods for longitudinal data are continually proposed to address complex real-world settings, simulation studies are essential to evaluate their effectiveness for practical use. These studies assess important properties such as bias, confidence interval coverage, relative efficiency, and robustness to violations of the assumptions used, guiding both methodological development and application. Sample size calculations also benefit from simulations. Prior to conducting a trial or an observational study, it is crucial to determine the number of participants or observed units required to achieve adequate statistical power. Although standard formulas can sometimes be used for time-fixed treatments, more complex scenarios with time-varying treatments make deriving analytical formulas impractical or impossible. In such cases, simulation is invaluable, allowing researchers to estimate sample sizes based on realistic data-generating processes and causal structures.
@@
A fundamental requirement for both method evaluation and sample size determination is the ability to simulate data that conform to a specified causal model. However, simulating longitudinal data from a pre-specified marginal structural model is not a straightforward task. While generating data sequentially from structural equations—creating covariates, treatment status, and then the outcome conditioned on past information—may seem intuitive, these models rely on conditional rather than marginal specifications. This makes it difficult to sequentially simulate from conditional distributions that align with a pre-specified marginal model; in particular, it is extremely challenging to sample from a null hypothesis—the so-called g-null paradox (Robins and Wasserman, 1997).
@@
The remainder of this paper is structured as follows: Section 2 reviews existing methods for simulating longitudinal data from causal models. Section 3 introduces key preliminaries, including the frugal parameterization (Evans and Didelez, 2024) and copulas. In Section 4, we introduce our proposed methodology. Section 5 extends the approach to time-to-event data, enabling simulations from survival models such as Cox marginal structural models, and Section 5.2 illustrates the method through a simulation study replicating a real-world clinical scenario involving concomitant treatment in a trial. We conclude with a discussion in Section 6. Example code for the simulations in Sections 4.4 and 5.2 is available at https://github.com/XiLinStats/LongSiMSM.
@@ 2 Related Literature
    Simulating from conditional models and deriving or approximating the corresponding marginal models post hoc.
@@
Several studies have explored the first approach. Xiao et al. (2010) used sequential conditional models to approximate Cox marginal structural models under the rare disease assumption, i.e., very low failure rates, which mitigates the impact of non-collapsibility but assumes the absence of unmeasured confounding between time-varying confounders and outcomes. Young and Tchetgen Tchetgen (2014) employed the g-formula to bridge conditional and marginal models in discrete-time Cox marginal structural models. While theoretically feasible for discrete time-varying covariates, deriving marginal models becomes intractable for continuous covariates due to the need for numerical integration. Simplifications, such as the Markov property or rare disease assumptions, are often necessary but limit generalizability. Sofrygin et al. (2017) developed the simcausal R package for flexible simulation from non-parametric structural equation models. The workflow involves specifying structural equations and sequentially simulating data from conditional distributions. However, deriving analytic expressions for marginal causal quantities under this conditional parameterization is often intractable and prone to model misspecification. To address this, users typically specify low-dimensional working marginal structural models (Neugebauer and van der Laan, 2007) and estimate their parameters via Monte Carlo simulations of counterfactual data. This introduces two limitations: first, the true marginal model remains unknown; second, reverse-engineering simulations to match a target causal function is challenging. Keogh et al. (2021) described simulation from additive-hazard marginal structural models (Aalen, 1989). They utilized the result that if the conditional hazard model follows an additive form, the equivalent marginal hazard model will also be additive. While the marginal model retains the correct form, true parameter values are generally intractable and require Monte Carlo approximations. A further limitation is the difficulty in ensuring non-negative hazards for continuous covariates.
@@
Alternative methods focus on simulating directly from the target marginal model. Havercroft and Didelez (2012) simulated directly from a discrete Cox marginal structural model, introducing time-varying confounding indirectly via a shared latent variable. However, this approach assumes no direct effect of time-varying confounders on outcomes, which may not hold in practice, e.g., the effect of CD4 count on HIV survival (Robins et al., 2000). Evans and Didelez (2024) proposed a frugal parameterization that prioritizes the causal margin and parameterizes the surrounding variables accordingly, enabling direct simulation from the exact marginal model while explicitly modeling time-varying covariates. This addresses key limitations of earlier methods. Building on this, Seaman and Keogh (2024) applied the frugal parameterization to simulate survival outcomes, using a one-dimensional “risk score” to summarize multi-dimensional covariates. While simplifying dependence structures, this method relies on computationally expensive Monte Carlo approximations to estimate the cumulative distribution functions of the risk score. In settings with continuous treatments or baseline variables, these approximations must be performed separately for each individual at each time point, substantially increasing the computational burden.
@@
The aims for a new simulation approach therefore include the ability to: simulate exactly from a specified marginal structural model; include multiple time-varying covariates, treatments, outcomes, and censoring; have good control over their distributions; avoid restrictive assumptions about the data-generating process. In addition, we would like it to scale well in the number of time-points and variables, and applicable in real-world settings.
@@ 3.2 Copulas
We can think of continuous multivariate distributions as the joining together of the univariate marginals (or margins) and the dependence structure. This structure can be represented by a copula, which is a multivariate distribution function $C:[0,1]^{d}\rightarrow[0,1]$ with margins uniformly distributed on $(0,1)$. According to Sklar’s theorem (Sklar, 1959, 1973), let $F$ be a multivariate distribution with margins $F_{1},\ldots,F_{d}$. Then there exists a copula $C$ such that for all $(x_{1},\ldots x_{d})$ in $\mathbb{R}^{d}$,
@@
where $u_{i}=F_{i}(x_{i})$, $i=1,\ldots,d$. If $F_{1}(x_{1}),\ldots,F_{d}(x_{d})$ are all continuous, then $C$ is unique and takes the form
@@
Otherwise, $C$ is only uniquely determined on $\operatorname{Ran}F_{1}\times\cdots\times\operatorname{Ran}F_{d}$, where $\operatorname{Ran}F_{i}$ denotes the range of $F_{i}$. The choice of copula can be considered independently of the univariate margins.
@@ 3.2.1 Pair-copula construction
Following the general formula given in Joe (1997), and assuming that the joint distribution of $(Y,X_{1},\ldots,X_{d})$ is absolutely continuous with continuous margins, the conditional distribution of $Y$ given $d$ other variables can be decomposed into $d$ bivariate copula density functions, each of which can be specified flexibly:
@@ 3.2.2 Simulating from pair-copula constructions
The pair-copula construction includes conditional distributions of the form $F(y\,|\,\bm{x})$. Assuming that the joint distribution of $(Y,X_{1},\ldots,X_{d})$ is absolutely continuous with continuous margins, and the bivariate copula $C_{YX_{j}\mid\bm{X}_{-j}}$ is differentiable, Joe (1996) demonstrated the following result:
$F(y\,|\,\bm{x})=\frac{\partial C_{YX_{j}\mid\bm{X}_{-j}}\left\{F(y\,|\,\bm{x}_{-j}),F(x_{j}\,|\,\bm{x}_{-j})\right\}}{\partial F(x_{j}\,|\,\bm{x}_{-j})},$ (3)
where $j\in\{1,\ldots,d\}$. For notational convenience, we introduce the h-function $h(u_{1},u_{2};\Theta)$ from Aas et al. (2009). Let $u_{1}$ and $u_{2}$ be random variables with a uniform distribution on $(0,1)$, for a sufficiently regular copula $C_{12}(u_{1},u_{2})$, the $h$-function is defined as
@@
where the second argument in $h(\cdot;\Theta)$ is the conditional variable, and $\Theta$ represents the set of copula parameters. The $h$-function is effectively the conditional copula $C(u_{1}\,|\,u_{2};\Theta)$. We further define $h^{-1}(u_{1},u_{2};\Theta)$, be the inverse of the $h$-function in the first variable $u_{1}$. Equivalently, $h^{-1}(u_{1},u_{2};\Theta):=F^{-1}(u_{1}\,|\,u_{2})$.
@@
Here we define a copula $C$ as sufficiently regular if it has non-zero density everywhere on its domain. Under this condition, the associated $h$-functions exist and are invertible. For standard parametric families—including the Gaussian, Student-t, Clayton and Frank copulas—the analytical form of $h(\cdot)$ and its inverse $h^{-1}(\cdot)$ can be derived directly from their definitions. See Appendix C of Aas et al. (2009).
@@ 4 Simulation Algorithm
We first consider a longitudinal study with $n$ individuals observed at discrete time points, $\mathcal{K}=\{0,\ldots,K\}$. At each follow-up $k$, treatment status $A_{k}$, and a set of time-varying covariates $L_{k}$ are recorded. Additionally, $Z$ represents a set of static baseline covariates. Latent variables, if present, are written as $\{H_{k}\}_{k\in\mathcal{K}}$. To indicate the history of a variable up to time $k$, we use bar notation. For instance, $\bar{A}_{k}=\{A_{0},\ldots,A_{k}\}$ and $\bar{L}_{k}=\{L_{0},\ldots,L_{k}\}$, with the convention $\bar{A}_{-1}=\emptyset$ and $\bar{L}_{-1}=\emptyset$. The outcome $Y$ is observed at the final time point $K$. Figure 1 is a graphical representation of this setup, which is common in medical and epidemiological research. Examples include examining the effects of an HIV treatment on infection status assessed at the end of the study period (Robins et al., 2000), and assessing the impact of iron supplement use during pregnancy on the odds of anemia at delivery (Bodnar et al., 2004). In Section 4.4, we illustrate our method through a simulation study that replicates such settings.
@@ 4.2 Frugal parameterization for outcome simulation
$F(y\,|\,\bar{l}_{K},\bar{a}_{K})$ $=h\big{[}\dots h\big{\{}h(F_{Y\mid\bar{A}_{K}}^{*}(y\,|\,do(\bar{a}_{K})),\,F(l_{0});\Theta_{0}),\,F(l_{1}\,|\,l_{0},a_{0});\Theta_{1}\big{\}}\dots;\Theta_{K}\big{]},$ (8)
which is centered around the causal margin $F^{*}_{Y\mid\bar{A}_{K}}$.
We write $h_{k}(u_{1},u_{2})=h(u_{1},u_{2};\Theta_{k})$, so as to avoid including the parameters $\Theta_{k}$. By the inversion method, it is straightforward that given ($\bar{l}_{K},\bar{a}_{K}$) and $U\sim\operatorname{U}(0,1)$,
$Y$ $=F^{*-1}_{Y\mid\bar{A}_{K}}(h_{0}^{-1}\big{[}\dots h_{K-1}^{-1}\big{\{}h_{K}^{-1}(U,F(l_{K}\,|\,\bar{l}_{K-1},\bar{a}_{K-1})),F(l_{K-1}\,|\,\bar{l}_{K-2},\bar{a}_{K-2})\big{\}},\ldots,F(l_{0})\big{]})$
follows the distribution parameterized by (7).
This result means that following the approach outline in Section 4.1, given the simulated baseline variables and time-varying variables, we can simulate the outcome in two steps:
@@
    $u^{*}_{Y}$ $=h_{0}^{-1}\big{[}\dots h_{K-1}^{-1}\big{\{}h_{K}^{-1}(w_{Y},F(l_{K}\,|\,\bar{l}_{K-1},\bar{a}_{K-1})),F(l_{K-1}\,|\,\bar{l}_{K-2},\bar{a}_{K-2})\big{\}}\dots,F(l_{0})\big{]}$ (9)
    $=h_{0}^{-1}\big{[}\dots h_{K-1}^{-1}\big{\{}h_{K}^{-1}(w_{Y},w_{K,1}),w_{K-1,1}\big{\}},\ldots,w_{0,1}\big{]},$
@@
An intuitive way to understand the calculation in Step 1 is that the inverse $h$-functions “unravel” the dependencies between $Y$ and each of $L_{K},L_{K-1},\dots L_{0}$ sequentially, and remove the conditioning of $Y$ on $\bar{L}_{K}$ until the marginal quantile of $Y$ is obtained. The quantiles in the inverse $h$-functions, such as $F(l_{k}\,|\,\bar{l}_{k-1},\bar{a}_{k-1})$, correspond exactly to the random standard uniform variables $w_{k,1}$ used to simulate $l_{k}$ at time step $k$. This enables direct substitution of the $l_{k}$ quantiles into (9) without additional computation.